# Grounding LLMs For Robot Task Planning Using Closed-loop State Feedback

### v2

@@ 3.1 Planning
Planning Prompt for Brain-LLM You are in the command of a mechanical robot. Your task is to split a given task into high-level steps that can be executed by the agent in the current environment. Each output step should be executable by the agent using available actions. Some examples of Task Instruction - Step-by-step plan pairs are given below: {in_context_learning_examples} You have the following objects in scene: {object_list}. The list of available actions are - {actions_available}. Use the information above to create the step-by-step plan for the given task instruction. Remember to only use the above objects and the available actions. Do not combine intermediate steps to generate compound steps. Make sure to complete all the steps needed to finish the task. Your reply should always start with ”0:” Environment Information: {environment_information} High-level Instruction: {task_input} Step-by-step Instructions:
@@
High level Plan BodyLLM Embedding Matching
[…]
@@ 3.3 Feedback
Error resolution in high level task planning has been explored through three primary sources of feedback: rule-based heuristics [14, 17], simulator feedback [34, 38, 43], and human feedback [30, 19]. Our method can integrate feedback from any of these sources to prompt the Brain-LLM to update its plans. The feedback prompt informs the Brain-LLM about the occurrence of an execution error and any associated error messages from the controller, simulator, or a human. Examples of error instances and corresponding solutions through new plans are also provided. Similar to the planning prompt, these in-context examples can be chosen either manually or from a suitable dataset. Each example consists of the tuple (Error Message, Explanation, Updated Plans). In our experiments, error messages are collected from the simulator or controller and indicate which action was not executed and the reason. Often, the raw error message is unreadable, so the Explanation explains the problem to the Brain-LLM in natural language and why a particular step failed. LLMs have been shown to improve performance with such chain-of-thought reasoning steps [50]. Finally, the updated plans are constrained to start from the error step and are conditioned on plans completed before the error step. In real-world applications, it is not feasible to restart the entire task since the robot might have already executed some plans. Thus, we constrain the Brain-LLM to always generate updated plans from the current error step. When deployed to resolve a particular error, the feedback prompt encourages the LLM to first generate a natural language reasoning for the failure, followed by step-by-step plans for error resolution. Figure 4 shows the prompt we use for Brain-LLM. In this work, we do not include any visual perception models. For simplicity, our method assumes that the locations of objects are known to the robot beforehand.
3.4 Differences from Previous Works
Our work derives inspiration from previous methods in utilizing feedback by back-prompting LLMs for improved robotic task planning [46]. A hierarchical planner, which breaks down complex task instructions to intermediate planning and action primitives has been explored before [42]. However, we design our system to ensure minimal domain customization, and demonstrate applicability in both simulation and real world setups. Rana et al. [34] uses two distinct LLMs for 3D scene graph search and iterative re-planning using simulator feedback. Our approach does not rely on scene graphs, with planning and error resolution performed by the Brain-LLM and action mapping performed by Body-LLM. Huang et al. [19] implements a closed loop feedback integrating visual perception and language conditioned robotic skills. In their approach, the LLM planner has access to vision models providing information about the environment, and it further disambiguate planning confusion by asking questions to a human operator. Our contributions lie in an effective and completely autonomous planning approach that eliminates any human-in-the-loop scenario by reinforcing feedback through error messages. There are a number of works that integrate visual grounding in planning with VLMs for object manipulation [48, 25, 23, 18] and navigation tasks [40, 53], and we leave such expansion of our planning methodology for future work. Our work can also be integrated with skill learning techniques through human-guided imitation learning as demonstrated in Parakh et al. [30].
A two-LLM system for generating symbolic and geometric relationships for task and motion planning has shown effectiveness in robotic environments [14]. Our method distinguishes itself by avoiding any human expert intervention or customization for filtering, allowing environmental feedback and error messages to automatically guide the planning LLM towards an accurate and correct real world action. Silver et al. [38] uses two LLMs for planning and correction using execution feedback, but relies on PDDL models for effective grounding, whereas our approach eliminates this dependency and can be used in real world experiments where objects and environment states are not known during task planning.
@@ 4.1 VirtualHome: Simulator For Embodied Control
VirtualHome (VH) serves as our robotic control software, simulating a Human-In-A-Household scenario with support for multiple agents [31]. The simulator features a variety of interactive household objects with predefined states like “open,” “closed,” “on,” and “off.” VH represents the agent as a humanoid avatar capable of interacting with these objects via low level control statements. Additionally, the simulator includes an in-built perception module that provides real-time information about objects in the scene, their states, and positions. Success and failure messages (with reasoning) for robotic commands are also provided. We use the environmental states and error messages from the in-built simulator as input to the feedback prompt described in Figure 4. For our experiments, we use the latest VirtualHome v2.3.0.
@@ 4.2 Dataset
We observed that many tasks from the original VH dataset, which were manually annotated, are no longer executable in the latest simulator version. Therefore, we utilize the train-validation-test subset of samples as used in Singh et al. [36]. Each sample comprises a tuple: a task description, a step-by-step plan (both in natural language), and corresponding VH commands. The dataset is divided into training, validation, and test splits, containing 35, 25, and 10 tuples, respectively. Our planning prompt uses examples of task descriptions and high-level plans from the dataset, whereas the execution prompt uses high-level plans and their corresponding VH commands as in-context learning examples.
@@ 4.3 Models Used
Since our approach relies on pre-trained foundational models that embody real-world knowledge and support incontext learning through prompting, we use three popular LLMs trained on vast amounts of data: PaLM 2 text-bison-001 [4], GPT-3.5 [9], and GPT-4 [1]. These models are accessed via their API calls.
@@ 4.4 Baselines
For fair comparisons, we selected baselines tested in similar scenarios—without visual perception-based grounding or human feedback. Our study benchmarks against two recent works that use LLMs for agent control in the VH environment [36, 17]. Additionally, we developed a baseline model (Baseline-LLM) to translate task descriptions, expressed in natural language, directly into VH commands for execution without any intermediate high-level plans. Unlike the BrainBody-LLM, the Baseline-LLM uses a single LLM without any environmental feedback but with the same training dataset. These comparisons will highlight the efficacy of our BrainBody-LLM in the context of LLM-based embodied control. Some in-context learning examples used in our prompt structure are described in Appendix A.
@@ 4.5 Evaluation Metrics
We use three metrics for evaluating the plans produced by BrainBody-LLM and the baselines.
@@ 4.6 Observations and Results
Table 2: Comparing LLM-based planning algorithms for the Virtual Home Simulator Environment
@@
Huang et al. [17] GPT-3 0.00$\pm$0.00 0.21$\pm$0.03
ProgPrompt[36] GPT-3 0.34$\pm$0.08 0.65$\pm$0.05
CODEX 0.40±$\pm 0.11$ 0.67±$\pm 0.08$
GPT-4 0.42±$\pm 0.08$ 0.65±$\pm 0.04$
Base-LLM PALM 0.30$\pm$0.00 0.38$\pm$0.08
GPT-3.5 0.16$\pm$0.05 0.68$\pm$0.03
BB-LLM PALM 0.30$\pm$0.08 0.61$\pm$0.13
GPT-3.5 0.26$\pm$0.15 0.57$\pm$0.06
GPT-4 0.38$\pm$0.13 0.70$\pm$0.01
BB-LLM PALM 0.40$\pm$0.00 0.59$\pm$0.01
GPT-3.5 0.36$\pm$0.18 0.72$\pm$0.04
GPT-4 0.54$\pm$0.09 0.69$\pm$0.04
@@
Obtained results demonstrate that feedback based planning and correction methods such as ProgPrompt and BrainBody-LLM (with feedback) outperform baselines that do not use feedback. Without feedback based error resolution, non-realizable plans are skipped which leads to unsatisfied goal conditions. ProgPrompt uses assert statements (if-else conditionals) to validate necessary environmental conditions before executing a given step. BrainBody-LLM (with feedback) uses error messages (in natural language) to guide the Brain-LLM towards a better plan and subsequent improved goal conditions satisfaction. Our approach exhibited overall enhancement in both SR and GCR metrics, outperforming ProgPrompt with the same LLM backend of GPT-4. This highlights the limitation of using simple assert based feedback signals. Our approach allows the LLM to comprehend raw error messages from the simulator, and with its advanced reasoning capabilities, self-diagnose and correct generated plans for improved success. Modifying the backend LLM from GPT-3.5 to GPT-4 improves our scores, motivating further improvements as LLMs become more powerful and their context window increases. In our approach, GPT-4 was the top performer, followed by PaLM 2 text-bison-001 and GPT-3.5.
@@
Figure 6: Impact of Feedback Loops on Evaluation Metrics. The graph demonstrates a clear upward trend in all evaluation metrics as the number of feedback loops increases, without significant fluctuations. Notably, upon incrementing the value of K beyond 3, the metrics stabilize around a consistent value. This value aligns with the results reported at K=3; therefore, data for values higher than K=3 are not included in the analysis.
@@ 5 Experiments with the Franka Robotic Arm
Figure 7: Demonstration of our BrainBody-LLM algorithm (GPT-4 - Backend), operating on a Franka Research 3 Robotic Arm. It successfully completes the sixth task in table 4 by creating a ”Plus” sign right side of the environment and subsequently returning to its initial position.
In this section, we demonstrate how BrainBody-LLM extends to real-world robot task planning using the Franka Research 3 (FR3) robotic arm. Our experiments, integrating a 7 DOF robotic arm with advanced LLMs like PaLM 2, GPT-3.5, and GPT-4, show promising results in autonomous robotic task planning and execution. All our experiments assume that the initial locations of objects and available target locations are known to the LLMs through environmental information communicated via prompts. This assumption can be removed with visual perception modules that identify and locate objects [28].
@@ 5.2 Results on Franka Research 3 Simulation Environment
We used a popular simulation environment for the FR3, which employs differential optimization and control barrier functions for pick-and-place tasks [12]. Scene objects were created with MuJoCo [44], and we developed scripts for API integration with three LLMs—PaLM 2, GPT-3.5, and GPT-4—using the prompt structure described in Section 3. Our tabletop environment consisted of objects distributed in both accessible and inaccessible parts of the workspace. This setup tested the LLM-based planner’s ability to adapt to out-of-workspace errors in real-time and revise plans accordingly. Extensive trajectory testing was conducted to avoid singularity points. We used a Python-based control repository [^1] https://github.com/Rooholla-KhorramBakht/FR3Py for the FR3 arm.
@@ 5.4 Guidelines for Adapting BrainBody-LLM to Diverse Robotic Environments
The BrainBody-LLM framework is designed to enable direct autonomous planning across a broad spectrum of robotic tasks, both in simulation and real-world environments. Apart from our experiments in VirtualHome, datasets such as ALFRED [41], VRKitchen [15], and TEACh [32] also employ high-level plans for task execution, making them ideal candidates for integration with our framework. Given that prompt-based experiments can be time-consuming, we propose a streamlined yet effective approach for adapting BrainBody-LLM to custom robotic tasks and environments:
@@ 6 Conclusion
This paper introduced an algorithm for robotic task planning that leverages the reasoning and dynamic error correction capabilities intrinsic to LLMs. Our approach improves both the success rate of task execution and the recall of goal conditions, positioning it favorably against current baselines within the Virtual Home Human-In-A-Household simulation environment. The design and conceptual framework of our algorithm are practical and readily adaptable for real-world implementation, as demonstrated through our experiments using the Franka Research 3 Robotic Arm in Section 5.
Future work will focus on further dissecting and mitigating the phenomena of closed-loop oscillations and hallucinations observed in LLM-generated plans. This will involve using multi-modal feedback through various sensor modalities to ensure grounded and realistic task planning. Our goal is to evaluate and refine the efficacy of LLM planning algorithms, thereby accelerating their adoption for robotics applications.
@@ Appendix A Prompt Examples for VirtualHome Experiments
[…]
@@ Appendix B Prompt Examples for Franka Arm Experiments
[…]
@@
[…]

### v3

@@ 3.1 Planning
Planning Prompt for Brain-LLM You are in the command of a mechanical robot. Your task is to split a given task into high-level steps that can be executed by the agent in the current environment. Each output step should be executable by the agent using available actions. Some examples of Task Instruction - Step-by-step plan pairs are given below: {in_context_learning_examples} You have the following objects in scene: {object_list}. The list of available actions are - {actions_available}. Use the information above to create the step-by-step plan for the given task instruction. Remember to only use the above objects and the available actions. Do not combine intermediate steps to generate compound steps. Make sure to complete all the steps needed to finish the task. Your reply should always start with “0:” Environment Information: {environment_information} High-level Instruction: {task_input} Step-by-step Instructions:
@@ 3.3 Feedback
Error resolution in high level task planning has been explored through three primary sources of feedback: rule-based heuristics [huang2022language, 10342169], simulator feedback [rana2023sayplan, Sun2023AdaPlannerAP, Silver_Dan_Srinivas_Tenenbaum_Kaelbling_Katz_2024], and human feedback [parakh2023humanassisted, pmlr-v205-huang23c]. Our method can integrate feedback from any of these sources to prompt the Brain-LLM to update its plans. The feedback prompt informs the Brain-LLM about the occurrence of an execution error and any associated error messages from the controller, simulator, or a human. Examples of error instances and corresponding solutions through new plans are also provided. Similar to the planning prompt, these in-context examples can be chosen either manually or from a suitable dataset. Each example consists of the tuple (Error Message, Explanation, Updated Plans). In our experiments, error messages are collected from the simulator or controller and indicate which action was not executed and the reason. Often, the raw error message is unreadable, so the Explanation explains the problem to the Brain-LLM in natural language and why a particular step failed. LLMs have been shown to improve performance with such chain-of-thought reasoning steps [NEURIPS2022_9d560961]. Finally, the updated plans are constrained to start from the error step and are conditioned on plans completed before the error step. In real-world applications, it is not feasible to restart the entire task since the robot might have already executed some plans. Thus, we constrain the Brain-LLM to always generate updated plans from the current error step. When deployed to resolve a particular error, the feedback prompt encourages the LLM to first generate a natural language reasoning for the failure, followed by step-by-step plans for error resolution. Figure 4 shows the prompt we use for Brain-LLM. In this work, we do not incorporate visual perception models. However, our framework is designed to be easily extendable to include such models, such as Grounding DINO [liu2023grounding], as visual perception modules to inform the BrainLLM about the environment. Integrating these models would enable continual feedback for reversible tasks. For instance, if a task involves pouring water into a glass but the water is accidentally spilled on the table, continual visual perception could provide feedback to the BrainLLM for re-planning. To simplify our approach, we assume that the robot has prior knowledge of object locations. This assumption is practical and can be relaxed using modern visual perception techniques.
@@ 4.1 VirtualHome: Simulator For Embodied Control
VirtualHome (VH) serves as our robotic control software, simulating a Human-In-A-Household scenario with support for multiple agents [puig_virtualhome]. The simulator features a variety of interactive household objects with predefined states like “open,” “closed,” “on,” and “off.” VH represents the agent as a humanoid avatar capable of interacting with these objects via low level control statements. Additionally, the simulator includes an in-built perception module that provides real-time information about objects in the scene, their states, and positions. Success and failure messages (with reasoning) for robotic commands are also provided. We use the environmental states and error messages from the in-built simulator as input to the feedback prompt described in Figure 4. For our experiments, we use VirtualHome v2.3.0.
@@ 4.2 Dataset
We observed that many tasks from the original VH dataset, which were manually annotated, are no longer executable in the latest simulator version. Therefore, we utilize the train-validation-test subset of samples as used in progprompt. Each sample comprises a tuple: a task description, a step-by-step plan (both in natural language), and corresponding VH commands. The dataset is divided into training, validation, and test splits, containing 35, 25, and 10 tuples, respectively. Our planning prompt uses examples of task descriptions and high-level plans from the dataset, whereas the execution prompt uses high-level plans and their corresponding VH commands as in-context learning examples.
@@ 4.3 Models Used
Since our approach relies on pre-trained foundational models that embody real-world knowledge and support incontext learning through prompting, we use three popular LLMs trained on vast amounts of data: PaLM 2 text-bison-001 [anil2023palm], GPT-3.5 [NEURIPS2020_1457c0d6], and GPT-4 [openai2023gpt4]. These models are accessed via their API calls. Furthermore, since Progprompt achieves optimal performance with the DAVINCI-003 model [NEURIPS2020_1457c0d6], a variant of GPT-3, we also report its performance using the DAVINCI-003 backbone.
@@ 4.4 Baselines
For fair comparisons, we selected baselines tested in similar scenarios—without visual perception-based grounding or human feedback. Our study benchmarks against two recent works that use LLMs for agent control in the VH environment [progprompt, huang2022language]. Additionally, we developed a baseline model (Baseline-LLM) to translate task descriptions, expressed in natural language, directly into VH commands for execution without any intermediate high-level plans. Unlike the BrainBody-LLM, the Baseline-LLM uses a single LLM without any environmental feedback but with the same training dataset. These comparisons will highlight the efficacy of our BrainBody-LLM in the context of LLM-based embodied control. Some in-context learning examples used in our prompt structure are described in Appendix A.
@@ 4.5 Evaluation Metrics
We use three metrics for evaluating the plans produced by BrainBody-LLM and the baselines as used in previous works [huang2022language, progprompt].
@@ 4.6 Observations and Results
Table 1: Comparing LLM-based planning algorithms for the Virtual Home Simulator Environment. We do not use the DAVINCI-003 LLM since it has been deprecated.
@@
[…]
DAVINCI-003 $\pm 0.15 0.74$$\pm$0.07
[…]
@@
Obtained results demonstrate that feedback based planning and correction methods such as ProgPrompt and BrainBody-LLM (with feedback) outperform baselines that do not use feedback. Without feedback based error resolution, non-realizable plans are skipped which leads to unsatisfied goal conditions. ProgPrompt uses assert statements (if-else conditionals) to validate necessary environmental conditions before executing a given step. BrainBody-LLM (with feedback) uses error messages (in natural language) to guide the Brain-LLM towards a better plan and subsequent improved goal conditions satisfaction. Our approach exhibited a significant enhancement in the SR metric, outperforming ProgPrompt with its highest-scoring LLM backend, Davinci-003, while achieving improved GCR when compared with the same LLM backend. Modifying our backend LLM from GPT-3.5 to GPT-4 improves our scores, motivating further improvements as LLMs become more powerful and their context window increases. In our approach, GPT-4 was the top performer, followed by PaLM 2 text-bison-001 and GPT-3.5. We also observe that BrainBodyLLM outperforms the Base-LLM baseline significantly on both SR and GCR when using the same LLM backend. Base-LLM was prompted to directly output low level control policies from task instructions. This shows that using two LLM calls for high level planning and low level policy generation leads to better results since LLMs are better at comprehending natural language prompts rather than functional APIs in their in-context learning examples. This motivates us to utilize a hierarchical structure for task planning, separating high-level task planning from low-level policy generation.
@@
Figure 6: Impact of Feedback Loops on Evaluation Metrics. The graph demonstrates a clear upward trend in all evaluation metrics as the number of feedback loops increases, without significant fluctuations. Notably, upon incrementing the value of K beyond 3, the metrics stabilize around a consistent value. This value aligns with the results reported at K=3; therefore, data for values higher than K=3 are not included in the analysis. While the graph shows results with GPT 3.5, similiar trend is observed with GPT 4.
High level Plan BodyLLM Embedding Matching
[…]
@@ 5 Experiments with the Franka Robotic Arm
Figure 7: Demonstration of our BrainBody-LLM algorithm (GPT-4 - Backend), operating on a Franka Research 3 Robotic Arm. It successfully completes the sixth task in table 4 by creating a “Plus” sign right side of the environment and subsequently returning to its initial position.
In this section, we demonstrate how BrainBody-LLM extends to real-world robot task planning using the Franka Research 3 (FR3) robotic arm. Our experiments, integrating a 7 DOF robotic arm with advanced LLMs like PaLM 2, GPT-3.5, and GPT-4, show promising results in autonomous robotic task planning and execution. All our experiments assume that the initial locations of objects and available target locations are known to the LLMs through environmental information communicated via prompts. This assumption can be removed with visual perception modules that identify and locate objects [minderer2022simple].
@@ 5.2 Results on Franka Research 3 Simulation Environment
We used a popular simulation environment for the FR3, which employs differential optimization and control barrier functions for pick-and-place tasks [Dai2023SafeNA]. Scene objects were created with MuJoCo [todorov2012mujoco], and we developed scripts for API integration with three LLMs—PaLM 2, GPT-3.5, and GPT-4—using the prompt structure described in Section 3. Our tabletop environment consisted of objects distributed in both accessible and inaccessible parts of the workspace. This setup tested the LLM-based planner’s ability to adapt to out-of-workspace errors in real-time and revise plans accordingly. Extensive trajectory testing was conducted to avoid singularity points. We used a Python-based control repository [^1] https://github.com/Rooholla-KhorramBakht/FR3Py for the FR3 arm.
@@ 5.4 Guidelines for Adapting BrainBody-LLM to Diverse Robotic Environments
The BrainBody-LLM framework is designed to enable direct autonomous planning across a broad spectrum of robotic tasks, both in simulation and real-world environments. Apart from our experiments in VirtualHome, datasets such as ALFRED [ALFRED20], VRKitchen [VRKitchen], and TEACh [Padmakumar2021TEAChTE] also employ high-level plans for task execution, making them ideal candidates for integration with our framework. Given that prompt-based experiments can be time-consuming, we propose a streamlined yet effective approach for adapting BrainBody-LLM to custom robotic tasks and environments:
@@ 6 Conclusion
This paper introduced an algorithm for robotic task planning that leverages the reasoning and dynamic error correction capabilities intrinsic to LLMs. Our approach improves the success rate of task execution and recall of goal conditions, positioning it favorably against current baselines within the Virtual Home Human-In-A-Household simulation environment. The design and conceptual framework of our algorithm are practical and readily adaptable for real-world implementation, as demonstrated through our experiments using the Franka Research 3 Robotic Arm in Section 5.
The assumption in our work regarding the reversibility of errors made by the robot introduces a notable limitation. The framework relies on the ability to create a new plan for achieving the given task after receiving environmental information and an error message. However, errors made by the planner might irreversibly alter the environment or damage the robot hardware, making the task unachievable. In future work, we aim to address this limitation by incorporating simulators into the planning loop and designing safety guarantees for the underlying action primitives used by Body-LLM. Our approach relies on implicit intervention through simulators or other forms of feedback communication with the environment. In the future, we aim to evaluate our approach in more complex and long-horizon scenarios compared to the experiments presented in this paper.
Future work will focus on understanding and mitigating the phenomena of closed-loop oscillations and hallucinations observed in LLM-generated plans. This effort will involve leveraging multi-modal feedback through diverse sensor modalities to ensure grounded and realistic task planning. Our overarching goal is to evaluate and refine the efficacy of LLM-based planning algorithms, thereby facilitating their broader adoption in robotics applications.
Acknowledgements This work was supported in part by ARO under Grant W911NF-22-1-0028 and in part by the New York University Abu Dhabi (NYUAD) Center for Artificial Intelligence and Robotics (CAIR), funded by Tamkeen under the NYUAD Research Institute Award CG010.
Author Contributions Conceptualization, V.B., A.K., P.K. and F.K.; methodology, V.B., A.K. and P.K.; software, V.B. and A.K.; resources, P.K., R.K., and F.K.; writing—original draft preparation, V.B. and A.K.; writing—review and editing, V.B., A.K., P.K., R.K. and F.K.; project administration, P.K., R.K. and F.K.; funding acquisition, P.K., R.K. and F.K.
Data availability No datasets were generated during the current study.
Competing interests The authors declare no competing interests.
Ethics Approval All authors consented to the submission of this manuscript and we explicitly state that the presented work is original and not submitted anywhere else. Our submission follows all ethical guidelines as outlined in the Autonomous Robots journal.
@@ Appendix A Prompt Examples for VirtualHome Experiments
You were given task: {input}, a generated Initial Plan was: {init_plan}. The robot received the following feedback message: {feedback_message} You have the following objects in scene: object list. The list of available actions are - walk, run, walktowards, walkforward, turnleft, turnright, sit, standup, grab, open, close, put, putin, switchon, switchoff, drink, touch and lookat. $\cdots$
@@ Appendix B Prompt Examples for Franka Arm Experiments
You have the following objects in scene: {object_list}. The list of available actions are - grasp, move, return to base. $\cdots$
@@
You were given task: {input}, a generated Initial Plan was: {init_plan}. The robot received the following feedback message: {feedback_message} You have the following objects in scene: object list. The list of available actions are - grasp, move, return to base. $\cdots$